A transformer has a voltage 480 V and

current 2.5 A on the primary side. And 80 V

on the secondary. What is the secondary

current on the load.


Select one:

a. 30 A

b. 20 A

C. 15 A

d. 10 A

e. 5 A


1
Expert's answer
2021-06-07T05:45:57-0400

"V_1=480V \newline
I_1 =2.5A\newline
V_2=80V \newline

We \space\frac{V_1}{V_2} =\frac{I_2}{I_1} \newline\newline

I_2 = \frac{400}{80}×2.5=15A"


Hence 15 A is the secondary current.